stylequeen:

Simplify square root of 3 multiplied by the fifth root of 3. 3 to the power of 1 over 10 3 to the power of 3 over 5 3 to the power of 9 over 10 3 to the power of 7 over 10

stylequeen:

can u help me

justus:

square root of 3 multiplied by the fifth root of 3 is (3)^1/2 * (3)^1/5 =

stylequeen:

huh

MrMudd183:

What do you think it is?

stylequeen:

I have no idea

MrMudd183:

First you would take the square root of 3, then after that you would multiply 3^5 `which is just 3*3*3*3*3`

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MrMudd183 What do you think it is? \(\color{#0cbb34}{\text{End of Quote}}\) This is a math question to solve, it's not about "what do you think it is".

MrMudd183:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @justus \(\color{#0cbb34}{\text{Originally Posted by}}\) @MrMudd183 What do you think it is? \(\color{#0cbb34}{\text{End of Quote}}\) This is a math question to solve, it's not about "what do you think it is". \(\color{#0cbb34}{\text{End of Quote}}\) I know, I was seeing what they thought the first step to do is or what the answer choice is

stylequeen:

243

MrMudd183:

That is one part of the of the answer, what do you think the next/final part is?

stylequeen:

i don't know srry

MrMudd183:

It's okay. So the way I would do it is to figure out what the square root of 3 is first

stylequeen:

ok

stylequeen:

1.73

MrMudd183:

Correct c: After that, I would multiply 3^5 `3*3*3*3*3`

stylequeen:

243

MrMudd183:

Correct. Now you would multiply those two together and then simplify it

jhonyy9:

@MrMudd183 sorry but i need asking what you know about the same numbers multiplied with different power what you need to do with exponents for example x^a *x^b = ?

MrMudd183:

I was taught that multiplying by an exponent is multiplying that number by itself the amount of times that the exponent is....such as 2^3 is 2*2*2

jhonyy9:

no this is very very complicated use the rule what i wrote above - and this is so easy

stylequeen:

I still have no idea what it is

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @justus square root of 3 multiplied by the fifth root of 3 is (3)^1/2 * (3)^1/5 = \(\color{#0cbb34}{\text{End of Quote}}\)

stylequeen:

I cant get the simplified version

jhonyy9:

what you know about this x^a *x^b = ?

stylequeen:

I know about it but I can't remember how to do this

jhonyy9:

really ? how you ve learned it ?

stylequeen:

I learned it in school but it has been a min

stylequeen:

I think it is the last one

jhonyy9:

ok so when you multiplie the same base numbers with different exponents write the base number and assume the exponents do you remember now ?